# Applying the Spectral Method

### v1

@@ 5. Testing Methodology and Computational Experiment
This paper concerns well-known linear filters with properties that are described in detail in various design manuals [1, 27, 28], so here testing does not mean an analysis of the amplitude and phase frequency characteristics of filters. Testing is related only to the spectral form, and it is based on a comparison of filter output signals with the original signal. This comparison is carried out in the spectral domain, and the transition to the time domain is implemented only for visualization of output signals.
@@
as the original signal and consider two noises. The first one is deterministic:
@@
The signal $g(t)=u(t)+v(t)$ is the filter input signal (an additive mixture of the original signal and noise), and the filter output $x(t)$ is compared with the original signal $u(t)$.
@@
The spectral characteristic $X=\mathbb{S}[x]$ of the output signal is given by the expression (5), in which the two-dimensional non-stationary transfer function $W$ is determined by one of the formulae from Section 4. The specific formula depends on the chosen filter family, filter order $n$, and cutoff frequency $\hat{\Omega}$. The output signal in the time domain is described by the formula (6).
For linear filters, even if there is no noise, i.e., $v(t)=0$, and the original signal corresponds to the passband, then we should not expect the equality $x(t)=u(t)$ to be satisfied. This follows from properties of solutions to linear differential equations (1). Therefore, the study of the error
@@
where $(\,\cdot\,)^{\scriptscriptstyle\mathrm{T}}$ means the transpose of a matrix. It is original to compare the error $\mathcal{E}$ with a priori error
@@
for the frequency $\omega$ that corresponds to the envelope of the original signal [27].
An important aspect of the spectral form of mathematical description is the necessity of truncating all infinite column matrices and infinite matrices to finite sizes for practical calculations. Without such a truncation, the formulae provided earlier remain of theoretical interest only. The truncation means that all column matrices for describing signals should have size $L$, and matrices for describing linear systems should have size $L\times L$. When transiting from the spectral domain to the time domain, the following relations are used for the output signal (without phase delay compensation and with phase delay compensation, respectively):
@@
A priori error $\mathcal{E}_{0}$ depends not only on $L$ but also on $\tau_{\varphi}$. This implies that it depends on the chosen filter family and the filter order $n$. For example, for the third-order Butterworth filter, a priori error has values $\mathcal{E}_{0}=0.242112,0.242318,0.242319,0.242319$ for $L=128,256,512,1024$, respectively. In order not to calculate a priori error for all considered filters, we restrict ourselves to the estimate for this characteristic:
@@
2 0.022445 0.022410 0.022303 0.022301
3 0.036950 0.036805 0.036705 0.036701
4 0.053139 0.052810 0.052770 0.052755
5 0.062381 0.061689 0.061610 0.061596
6 0.079740 0.078643 0.078510 0.078492
@@
Figure 5. The graphs of original signal and sum of original signal and deterministic noise
Selectively, i.e., for some $n$ and $L$, the filtering results are presented in Figures 99: for $u(t)$ (red color) and $x^{*}(t)$ (blue color). Noticeable discrepancies in the graphs of functions $g(t)$ and $x^{*}(t)$ occur only on $T-\tau_{\varphi},T]$, and this is quite expected. The data in tables indicate the effectiveness of deterministic noise suppression, since the errors $\mathcal{E}$ are several times smaller than values $\mathcal{E}_{0}^{+}$ which estimate a priori error.
Figure 6. The graphs of the original signal and the output signal of the Butterworth filter ($n=3$, $L=128$, deterministic noise)
Figure 7. The graphs of the original signal and the output signal of the Linkwitz–Riley filter ($n=4$, $L=256$, deterministic noise)
Figure 8. The graphs of the original signal and the output signal of the Chebyshev Type I filter ($n=5$, $L=512$, deterministic noise)
Figure 9. The graphs of the original signal and the output signal of the Chebyshev Type II filter ($n=2$, $L=1024$, deterministic noise)
@@
2 0.048527 (0.006313) 0.048451 (0.006310) 0.048447 (0.006290)
3 0.070567 (0.008081) 0.070621 (0.008134) 0.070434 (0.008009)
4 0.096689 (0.009690) 0.096749 (0.009839) 0.096935 (0.009852)
5 0.127415 (0.013328) 0.127424 (0.013173) 0.127248 (0.013351)
6 0.181272 (0.021788) 0.181213 (0.021586) 0.181538 (0.022092)
@@
Figure 10. The graphs of original signal and the sum of original signal and random noise
Figures 1414 show the results in graphical form for some values $n$ and $L$: for $u(t)$ (red color) and $x^{*}(t)$ (blue color); the graphs are shown for a single realization of $v(t)$. The filters are less effective here, but this is clear since the noise contains components with frequencies from the passband. Here, the filtering is quite effective except for the high-order Chebyshev Type II filters.
Figure 11. The graphs of the original signal and the output signal of the Butterworth filter ($n=3$, $L=128$, random noise)
Figure 12. The graphs of the original signal and the output signal of the Linkwitz–Riley filter ($n=4$, $L=256$, random noise)
Figure 13. The graphs of the original signal and the output signal of the Chebyshev Type I filter ($n=5$, $L=512$, random noise)
Figure 14. The graphs of the original signal and the output signal of the Chebyshev Type II filter ($n=2$, $L=1024$, random noise)
@@ 6. Conclusions
In the proposed technique, input and output signals are represented by non-stationary spectral characteristics formed by expansion coefficients of these signals relative to the basis of the space of square-integrable functions. Algebraic operations are used to transform non-stationary spectral characteristics. For testing, cosines are chosen as the basis, but the basis can be arbitrary.
The advantage of the considered technique is that the output signal is modeled in continuous time. In addition to the above filters, it can be used for elliptic filters, as well as for the Bessel, Gaussian, and Legendre filters, etc.
@@
Next, we consider how the redefinition of the matrix $S$ affects the error $\mathcal{E}$. In this case, we restrict ourselves to the Butterworth filters only since the conclusions for other filters are similar. The calculation results are given in Table 10 for deterministic noise and in Table 10 for random noise. The comparison of Tables 4 and 10, as well as 8 and 10 shows that the quality of modeling linear filters improves by increasing the filter order when the basis functions are defined beyond the segment $\mathds{T}$ in a natural way.
@@
Figure 15. The graphs of the original signal and the output signal of the Butterworth filter ($n=3$, $L=128$, deterministic noise, another definition of $S$)
Figure 16. The graphs of the original signal and the output signal of the Butterworth filter ($n=3$, $L=128$, random noise, another definition of $S$)

### v2

@@ 5. Testing Methodology and Computational Experiment
This paper concerns well-known linear filters with properties that are described in detail in various design manuals [1, 2, 27, 29], so here testing does not mean an analysis of the amplitude and phase frequency characteristics of filters. Testing is related only to the spectral form, and it is based on a comparison of filter output signals with the original useful signal. This comparison is carried out in the spectral domain, and the transition to the time domain is implemented only for visualization of output signals.
@@
as the useful signal and consider two noises. The first one is deterministic:
@@
The signal $g(t)=u(t)+v(t)$ is the filter input signal (the additive mixture of the useful signal and noise), and the filter output $x(t)$ is compared with the useful signal $u(t)$.
@@
The spectral characteristic $X=\mathbb{S}[x]$ of the output signal is given by the expression (5), in which the two-dimensional non-stationary transfer function $W$ is determined by one of the formulae from Section 4. The specific formula depends on the chosen filter family, filter order $n$, and cutoff frequency $\hat{\Omega}$. Further, we will assume that $\hat{\Omega}=40\pi$. The output signal in the time domain is described by the formula (6).
For linear filters, even if there is no noise, i.e., $v(t)=0$, and the useful signal corresponds to the passband, then we should not expect the equality $x(t)=u(t)$ to be satisfied. This follows from properties of solutions to linear differential equations (1). Therefore, the study of the error
@@
where $(\,\cdot\,)^{\scriptscriptstyle\mathrm{T}}$ means the transpose of a matrix. It is useful to compare the error $\mathcal{E}$ with a priori error
@@
for the frequency $\omega$ that corresponds to the envelope of the useful signal [27].
The spectral form of mathematical description has an important aspect: it requires truncating all infinite column matrices and infinite matrices to finite sizes when used in computer modeling. Without such a truncation, the formulae provided earlier remain of theoretical interest only. The truncation means that all column matrices for describing signals should have size $L$, and matrices for describing linear systems should have size $L\times L$. When transiting from the spectral domain to the time domain, the following relations are used for the output signal (without phase delay compensation and with phase delay compensation, respectively):
@@
The choice of truncation orders is motivated by several factors. The representation of the input signal as a partial sum of the Fourier series exhibits properties of a low-pass filter. To avoid this from affecting the results when testing linear filters, the truncation order should be chosen such that the additive mixture of the useful signal and noise is represented by the truncated non-stationary spectral characteristic without significant loss of accuracy. The truncation order $L=128$ satisfies this condition.
Furthermore, $\log_{2}128$ is an integer number. If only cosines are used as a basis, then this condition is not significant. However, such a choice is important for the Walsh and Haar functions (they are defined so that optimal truncation orders for them are $2^{\mu}$ for integer $\mu$). The truncation order $L=128$ for cosines allows future comparisons of the presented results with those obtained by the Walsh and Haar functions.
The truncation orders $L=256,512,1024$ are needed to analyze the dependence of the error on $L$ (doubling the truncation order is a common practice for the spectral method [33], and it is similar to halving the integration step for numerical methods).
A priori error $\mathcal{E}_{0}$ depends not only on $L$ but also on $\tau_{\varphi}$. This implies that it depends on the chosen filter family and the filter order $n$. For example, for the third-order Butterworth filter, a priori error has values $\mathcal{E}_{0}=0.242112,0.242318,0.242319,0.242319$ for $L=128,256,512,1024$, respectively. In order not to calculate a priori error for all considered filters, we restrict ourselves to the its estimate:
@@
2 0.040065 0.040035 0.039972 0.039971
3 0.026718 0.026629 0.026509 0.026503
4 0.014455 0.014213 0.014108 0.014072
5 0.022295 0.022069 0.022013 0.021994
6 0.022795 0.022501 0.022454 0.022444
Two components of the error $\mathcal{E}$ can be distinguished. The first one is the filter error, which is definitely present, since none of the filters under consideration are ideal. The second one is the spectral method error, associated with the truncation of all non-stationary spectral characteristics and two-dimensional non-stationary transfer functions. This component decreases with increasing truncation order $L$, which should be especially noticeable for smooth signals [40]. However, the errors presented in Tables 44 decrease rather slowly. These results indicate that the spectral method error is insignificant compared to the filter error.
The data in tables indicate the effectiveness of deterministic noise suppression, since the errors $\mathcal{E}$ are several times smaller than values $\mathcal{E}_{0}^{+}$ which estimate a priori error. In the considered example, an increase in $L$ is not required, since it does not lead to a noticeable improvement in the filter quality. The results from Tables 44 show that filter quality generally improves with increasing filter order.
@@
Figure 5. The graphs of the useful signal and the sum of useful signal and deterministic noise
Selectively, i.e., for some $n$ and $L$, the filtering results are presented in Figures 99: for $u(t)$ (red color) and $x^{*}(t)$ (blue color). Noticeable discrepancies in the graphs of functions $g(t)$ and $x^{*}(t)$ occur only on $T-\tau_{\varphi},T]$, and this is quite expected.
Figure 6. The graphs of the useful signal and the output signal of the Butterworth filter ($n=3$, $L=128$, deterministic noise)
Figure 7. The graphs of the useful signal and the output signal of the Linkwitz–Riley filter ($n=4$, $L=256$, deterministic noise)
Figure 8. The graphs of the useful signal and the output signal of the Chebyshev Type I filter ($n=5$, $L=512$, deterministic noise)
Figure 9. The graphs of the useful signal and the output signal of the Chebyshev Type II filter ($n=2$, $L=1024$, deterministic noise)
@@
2 0.055214 (0.007209) 0.056249 (0.007106) 0.058224 (0.006821)
3 0.052184 (0.006383) 0.052560 (0.006359) 0.052775 (0.006335)
4 0.056234 (0.006720) 0.056485 (0.006577) 0.058290 (0.006415)
5 0.058212 (0.006699) 0.059000 (0.006695) 0.059625 (0.006630)
6 0.060050 (0.006831) 0.059860 (0.006811) 0.061285 (0.006688)
The analysis of errors with statistical characteristics from Tables 88 enables the formulation of the same conclusion: the spectral method error is insignificant compared to the filter error. The filters are less effective here, but this is clear since the noise contains components with frequencies from the passband. These factors mean that both the truncation order and the filter order have no noticeable effect on the results. However, filters under random noise are also effective because the estimate of the mathematical expectation $\bar{\mathcal{E}}$ is less than the estimate of the mathematical expectation $\bar{\mathcal{E}}_{0}^{+}$.
@@
Figure 10. The graphs of the useful signal and the sum of useful signal and random noise
Figures 1414 show the results in graphical form for some values $n$ and $L=256$: for $u(t)$ (red color) and $x^{*}(t)$ (blue color); the graphs are shown for a single realization of $v(t)$.
Figure 11. The graphs of the useful signal and the output signal of the Butterworth filter ($n=3$, $L=256$, random noise)
Figure 12. The graphs of the useful signal and the output signal of the Linkwitz–Riley filter ($n=4$, $L=256$, random noise)
Figure 13. The graphs of the useful signal and the output signal of the Chebyshev Type I filter ($n=5$, $L=256$, random noise)
Figure 14. The graphs of the useful signal and the output signal of the Chebyshev Type II filter ($n=2$, $L=256$, random noise)
The computational complexity of finding the two-dimensional non-stationary transfer function for a linear filter can be estimated as $O(L^{3})$. This follows from properties of matrix multiplication and inversion, and this estimate holds when using standard algorithms. When using fast algorithms, the computational complexity is lower. For example, if the Strassen algorithm is applied, then the computational complexity can be estimated as $O(L^{\log_{2}7})$. The computational effort required to obtain the output signal, provided the known two-dimensional non-stationary transfer function, is proportional to the truncation order $L$.
@@ 6. Conclusions
In the proposed technique, input and output signals are represented by non-stationary spectral characteristics formed by expansion coefficients of these signals relative to the basis of the space of square-integrable functions. Algebraic operations are utilized to transform non-stationary spectral characteristics. For testing, cosines are chosen as the basis, but the basis can be arbitrary. In fact, the spectral method combines the advantages of analytical methods and the simple implementation inherent in numerical methods.
The advantage of the considered technique is that the output signal is modeled in continuous time. In addition to the above filters, it can be used for elliptic filters, as well as for the Bessel, Gaussian, and Legendre filters, etc. For these filters, the algorithm for deriving the transfer function is more complex. In particular, elliptic filters require the use of special functions (Jacobi elliptic functions), while the Gaussian filter is defined by an approximate representation of the Gaussian function as a partial sum of the Maclaurin series. However, once the transfer function is obtained, applying the spectral method for modeling these filters is no more difficult than for the Butterworth, Linkwitz–Riley, and Chebyshev filters.
@@
Next, we consider how the redefinition of the matrix $S$ affects the error $\mathcal{E}$. In this case, we restrict ourselves to the Butterworth filters only since the conclusions for other filters are similar. The calculation results are given in Table 10 for deterministic noise and in Table 10 for random noise. The comparison of Tables 4 and 10, as well as 8 and 10 shows that the filter quality improves by increasing the filter order when the basis functions are defined beyond the segment $\mathds{T}$ in a natural way.
@@
Figure 15. The graphs of the useful signal and the output signal of the Butterworth filter ($n=3$, $L=128$, deterministic noise, another definition of $S$)
Figure 16. The graphs of the useful signal and the output signal of the Butterworth filter ($n=3$, $L=128$, random noise, another definition of $S$)